# DynamicTree: Interactive Real Tree Animation via Sparse Voxel Spectrum

### v1

@@ 3.3 Voxel Grid Condition
To reduce the synthetic-to-real gap when using multi-view images as input, we condition the motion generation model on voxel grids. Given multi-view images of a static tree, we first reconstruct its mesh $M=(V,F)$ using an off-the-shelf method Guédon & Lepetit (2024), where $V=\{v_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ and $F=\{f_{j}\subset\{1,\dots,N\}\}_{j=1}^{P}$ denote the sets of vertices and faces, respectively. We then voxelize the mesh to obtain a sparse voxel grid $G$, which serves as the conditioning input.
@@ 3.4 Sparse Voxel Diffusion
Before performing the diffusion generation, the sparse voxel grid $G$ is encoded via a sparse encoder with several sparse convolutional blocks Williams et al. (2024), resulting in a compact latent representation $g\in\mathbb{R}^{d\times n}$ as geometric conditioning. Our sparse voxel diffusion module builds on the U-Net architecture introduced by XCube Ren et al. (2024). Specifically, to generate the sparse voxel spectrum $\mathcal{S}=\{s_{i}\in\mathbb{R}^{6\times n}\mid i=1,\dots,K\}$ of the mesh motion, we condition the diffusion generation process on both the frequency index and the latent feature $g$, generating each frequency component separately. The diffusion process Ho et al. (2020) starts from pure Gaussian noise and iteratively predicts noise over $L$ Markov steps. At each step, we concatenate the latent feature $g$ with the noisy latent $s_{l}$, and inject the frequency embedding into every ResBlock of the sparse voxel U-Net through scale and shift operations.
@@ 3.5 Optimization
where $g$ and $f$ denote the sparse voxel latent condition and frequency embedding, respectively. Further, we find that using only the diffusion loss $\mathcal{L}_{DM}$ may lead to unrealistic motion, such as divergence of some branches, as shown in Fig. 5, because the problem is under-constrained. To address this, we introduce a Local Spectrum Smoothness (LSS) loss that encourages local consistency in the frequency domain, inspired by the physical prior proposed by Sorkine & Alexa (2007) that neighboring points tend to move similarly. Specifically, we compute discrepancies in the frequency-domain parameters between each point and its neighbors, weighted by spatial proximity:
$\mathcal{L}_{\text{LSS}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j\in\mathcal{N}(i)}e^{-\alpha d_{ij}}\left(\|\text{Re}_{i}-\text{Re}_{j}\|+\lambda\|\text{Im}_{i}-\text{Im}_{j}\|\right),$
@@ 3.6 Mesh-Driven 3DGS Animation
With the recovered mesh motion, we then animate the 3DGS model, by binding Gaussian primitives to the mesh surface, as proposed by GaMeS Waczyńska et al. (2024). This operation can be viewed as a reparameterization: for each face $f_{j}=\{v_{1},v_{2},v_{3}\}\in\mathbb{R}^{3}$, we parameterize the attributes of its associated Gaussian primitive ($u$, $r$, and $s$) using the positions of the three vertices:
@@
where $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are learnable parameters, and $r_{1}$, $r_{2}$, $r_{3}$, $s_{1}$, $s_{2}$, and $s_{3}$ are parameterization functions. For details, please refer to Waczyńska et al. (2024). Through this binding strategy, we can compute the 3DGS deformation $\mathcal{D}_{g}$ directly from the mesh motion $\mathcal{D}_{m}$. This allows us to obtain the final deformed 3DGS sequence $\mathcal{G}$.
@@ 4 Interactive Simulation with Modal Analysis
Modal analysis is a technique used to decompose complex deformable motions into a set of fundamental vibration modes, each associated with a specific natural frequency. This approach is particularly well-suited for modeling the motion of systems composed of superpositions of harmonic oscillators, such as tree motion Diener et al. (2009); Habel et al. (2009). Given an external force $\mathbf{f}(t)$, we model all vertices of the tree mesh as a interconnected mass-spring-damper system $P$ to simulate the response $\mathcal{D}(t)=\{d_{i}(t)\in\mathbb{R}^{3}\mid i\in P\}$. With this, we can construct the following equation of motion Shabana (1991):
@@
To solve this equation, we project it into modal space, resulting in $|P|$ independent equations Davis et al. (2015); Li et al. (2024):
@@
where $m_{i}$, $c_{i}$, and $k_{i}$ correspond to the diagonal elements of the respective matrices. This is a standard second-order differential equation, which can be solved using explicit Euler integration. To perform the integration, we need to specify the initial modal displacement $q_{i}(0)$ and velocity $\dot{q}_{i}(0)$. These settings, along with the selection of M{M}, C{C}, and K{K}, are based on the configurations described in Petitjean et al. (2023).
@@
Thanks to prior work Davis et al. (2015); Li et al. (2024); Petitjean et al. (2023) that has shown the spectrums of particle motion trajectories can be treated as modal bases, we can use the mesh motion spectrums computed in Sec. 3 as the modal shapes $\phi$ to solve the above equation and obtain the interactive dynamic response $\mathcal{D}(t)$.
Figure 3: Comparison with 4DGen Yin et al. (2023). We visualize the middle frame of the generated sequence, where our method preserves better 3D structures. Space-time slices are shown, with vertical and horizontal axes representing time and the spatial profile along the brown line, respectively.
@@ 5 Dataset
To facilitate the learning of complex 3D tree dynamics, we introduce 4DTree, a large-scale 4D tree dataset containing $8,786$ animated tree meshes. Each instance includes a 100-frame animation, along with semantic labels for leaves and trunks, which can further support downstream tasks such as semantic segmentation.
To create 4D tree data, a straightforward approach is to use commercial physics-based simulation software, but this is time-consuming and impractical for large-scale datasets. Instead, we adopt the method from Weber & Penn (1995), which models trees as hierarchical branching structures and simulates wind-induced motion by treating stems as elastic rods coupled through oscillators, implemented via the Sapling Tree Gen add-on in Blender. Still, this approach involves many sensitive parameters, which, if set improperly, can cause unstable oscillations or unrealistic shapes. To ensure quality and consistency of our dataset, we adopt a three-stage pipeline during production: parameter tuning, automatic validation with scripts, and final visual filtering by human reviewers. Through this process, we construct a clean and diverse 4D tree dataset with complex dynamics. For Detailed procedures and data samples, please refer to the Appendix.
@@ 6 Experiments
Implementation. We train our model from scratch without pre-trained models, taking 3.5 days on 8 RTX 4090 GPUs with a batch size of 48. During the first 40,000 iterations, we train the model using only the $\mathcal{L}_{DM}$ loss. Then, we introduce the $\mathcal{L}_{LSS}$ loss and continue training for an additional 30,000 iterations. For the $\mathcal{L}_{LSS}$ loss, we use the 5 nearest neighbors of each point, and both $\alpha$ and $\lambda$ are set to 0.5. We set the resolution of the sparse voxel spectrum to $128^{3}$, with an input resolution of $512^{3}$ for the sparse voxel encoder. This results in voxel latent conditions of dimensionality $d=128$ at the $128^{3}$ resolution. When binding 3DGS primitives, we assign five Gaussians per face. We use the AdamW optimizer with an initial learning rate of $1\times 10^{-4}$, which is halved every 20,000 iterations. During inference, we employ DDIM Song et al. (2020) with 100 sampling steps.
Evaluation metrics. To evaluate our method in real-world scenarios, we collect a test set of 13 real-world trees. For evaluation metrics, we follow prior works Yin et al. (2023); Chen et al. (2024) and use CLIP ViT-B/32 Radford et al. (2021) to measure both visual realism and temporal coherence. Specifically, we compute CLIP-I distance as the average CLIP distance between each frame and the input view, and CLIP-T distance as the average CLIP distance between consecutive frames. Furthermore, we conduct a user study on the rendered videos, focusing on four key aspects: motion authenticity (MA), motion complexity (MC), 3D structural consistency (3DSC), and visual quality (VQ). Below, we compare the results of motion generation and interaction simulation, respectively.
@@ 6.1 Comparison of 3D Animation
We select 4DGen Yin et al. (2023) as the baseline for motion generation. As shown in Fig. 3, the results of 4DGen often exhibit artifacts in fine details. Moreover, when the tree structures become more complex, 4DGen would fail to converge due to the degradation of motion generation in its underlying VDM. For quantitative comparison, we evaluate only the input-view results to ensure a fair evaluation, as 4DGen performs poorly on novel views. The results in Table 1 demonstrate that our method achieves superior performance across all metrics. More animation results are reported in the Appendix. We strongly recommend viewing the supplementary videos for better visual assessment.
@@ 6.2 Comparison of Interactive Simulation
For interactive dynamic simulation, we compare against state-of-the-art baselines PhysGaussian Xie et al. (2024) and PhysFlow Comas et al. (2024). As shown in Fig. 4, PhysGaussian and PhysFlow often produce unrealistic plastic deformations. Specifically, PhysGaussian deforms slowly with little rebound, while PhysFlow exhibits partial recovery but still lacks fine-grained elasticity at the branch and leaf level, producing overly global responses. In contrast, our method produces natural and elastic motions, with branches and leaves exhibiting distinct behaviors. Quantitative results from four viewpoints and simulation time are reported in Table 1. Our method not only outperforms the baselines on most metrics but also significantly reduces the simulation time. For each frame, our method takes only about 18 ms for simulation, with 13 ms for mesh motion computation via modal analysis, 2.57 ms for Gaussian deformations calculation, and 2.65 ms for rendering, achieving a real-time interaction. Note that PhysFlow requires additional parameter optimization, which results in significantly longer runtime. More simulation results can be found in the Appendix.
@@ 7 Limitation and Conclusion
In this paper, we present DynamicTree, a novel framework for animating 3DGS trees. By introducing the sparse voxel spectrum representation, our method enables efficient long-term motion generation and real-time dynamic response to external forces. Furthermore, we also introduce a large-scale synthetic 4D tree dataset to support learning-based tree motion generation. Experimental results demonstrate that our approach achieves high-quality tree motion with strong temporal coherence and physical plausibility.
@@
Appendix A Appendix
A.1 Dataset
As discussed in the paper, we implement the method of Weber & Penn (1995) as a plugin in Blender. However, generating a large-scale 4D tree dataset using this method remains challenging, as it involves numerous parameters and random sampling often produces invalid or unrealistic 4D trees. To ensure data quality and consistency, we adopt a three-stage pipeline to construct our dataset:
@@ A.1 Dataset
Figure 6: Examples from our synthetic dataset. To demonstrate the semantic labels, we render the leaves and trunk with two simple material settings. Users can replace these with more realistic materials for enhanced visual quality.
A.2 Network
The sparse encoder and sparse voxel diffusion U-Net are adapted from the basic modules proposed in XCube to better fit our conditioning input and spectral output. We report key parameter settings of these two components in Table 3
@@ A.2 Network
A.3 Visualization results
Further visualizations and analytical details of the 3D animations and interactive dynamic simulations are presented in Fig. 8 and Fig. 7. However, to facilitate a comprehensive perceptual and qualitative evaluation of our method, we strongly recommend reviewing our supplementary videos.
@@ A.3 Visualization results
Figure 8: More results of 3D animation. For each scene, we visualize the space-time slices of depth and RGB videos from two viewpoints. We also show the scene flow of three mesh point cloud frames $(t_{1}=30,t_{2}=50,t_{3}=80)$ in the generated sequence, with color coding following the strategy used in Hur & Roth (2020), where the movements $(u,v)$ along the $x$ and $z$ directions are encoded using standard optical flow coloring.

### v2

@@ 3.3 Voxel Grid Condition
To reduce the synthetic-to-real gap when using multi-view images as input, we condition the motion generation model on voxel grids. Given multi-view images of a static tree, we first reconstruct its mesh $M=(V,F)$ using an off-the-shelf method [13], where $V=\{v_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ and $F=\{f_{j}\subset\{1,\dots,N\}\}_{j=1}^{P}$ denote the sets of vertices and faces, respectively. We then voxelize the mesh to obtain a sparse voxel grid $G$, which serves as the conditioning input.
@@ 3.4 Sparse Voxel Diffusion
Before performing the diffusion generation, the sparse voxel grid $G$ is encoded via a sparse encoder with several sparse convolutional blocks [46], resulting in a compact latent representation $g\in\mathbb{R}^{d\times n}$ as geometric conditioning. Our sparse voxel diffusion module builds on the U-Net architecture introduced by XCube [35]. Specifically, to generate the sparse voxel spectrum $\mathcal{S}=\{s_{i}\in\mathbb{R}^{6\times n}\mid i=1,\dots,K\}$ of the mesh motion, we condition the diffusion generation process on both the frequency index and the latent feature $g$, generating each frequency component separately. The diffusion process [15] starts from pure Gaussian noise and iteratively predicts noise over $L$ Markov steps. At each step, we concatenate the latent feature $g$ with the noisy latent $s_{l}$, and inject the frequency embedding into every ResBlock of the sparse voxel U-Net through scale and shift operations.
@@ 3.5 Optimization
where $g$ and $f$ denote the sparse voxel latent condition and frequency embedding, respectively. Further, we find that using only the diffusion loss $\mathcal{L}_{DM}$ may lead to unrealistic motion, such as divergence of some branches, as shown in Fig. 5, because the problem is under-constrained. To address this, we introduce a Local Spectrum Smoothness (LSS) loss that encourages local consistency in the frequency domain, inspired by the physical prior proposed by Sorkine and Alexa [40] that neighboring points tend to move similarly. Specifically, we compute discrepancies in the frequency-domain parameters between each point and its neighbors, weighted by spatial proximity:
$\mathcal{L}_{\text{LSS}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j\in\mathcal{N}(i)}e^{-\alpha d_{ij}}\left(\|\text{Re}_{i}-\text{Re}_{j}\|+\lambda\|\text{Im}_{i}-\text{Im}_{j}\|\right)$ (3)
@@ 3.6 Mesh-Driven 3DGS Animation
With the recovered mesh motion, we then animate the 3DGS model, by binding Gaussian primitives to the mesh surface, as proposed by GaMeS [43]. This operation can be viewed as a reparameterization: for each face $f_{j}=\{v_{1},v_{2},v_{3}\}\in\mathbb{R}^{3}$, we parameterize the attributes of its associated Gaussian primitive ($u$, $r$, and $s$) using the positions of the three vertices:
@@
where $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are learnable parameters, and $r_{1}$, $r_{2}$, $r_{3}$, $s_{1}$, $s_{2}$, and $s_{3}$ are parameterization functions. For details, please refer to [43]. Through this binding strategy, we can compute the 3DGS deformation $\mathcal{D}_{g}$ directly from the mesh motion $\mathcal{D}_{m}$. This allows us to obtain the final deformed 3DGS sequence $\mathcal{G}$.
@@ 4 Interactive Simulation with Modal Analysis
Modal analysis is a technique used to decompose complex deformable motions into a set of fundamental vibration modes, each associated with a specific natural frequency. This approach is particularly well-suited for modeling the motion of systems composed of superpositions of harmonic oscillators, such as tree motion [10, 14]. Given an external force $\mathbf{f}(t)$, we model all vertices of the tree mesh as an interconnected mass-spring-damper system $P$ to simulate the response $\mathcal{D}(t)=\{d_{i}(t)\in\mathbb{R}^{3}\mid i\in P\}$. With this, we can construct the following equation of motion [37]:
@@
Figure 3: Comparison with 4D generation methods. We visualize the middle frame of the sequence, where our method better preserves 3D structures. Space-time slices are shown, with vertical and horizontal axes representing time and the spatial profile along the brown line.
To solve this equation, we project it into modal space, resulting in $|P|$ independent equations [8, 23]:
@@
where $m_{i}$, $c_{i}$, and $k_{i}$ correspond to the diagonal elements of the respective matrices. This is a standard second-order differential equation, which can be solved using explicit Euler integration. To perform the integration, we need to specify the initial modal displacement $q_{i}(0)$ and velocity $\dot{q}_{i}(0)$. These settings, along with the selection of M{M}, C{C}, and K{K}, are based on the configurations described in [29].
@@
Thanks to prior work [8, 23, 29] that has shown the spectrums of particle motion trajectories can be treated as modal bases, we can use the mesh motion spectrums computed in Sec. 3 as the modal shapes $\phi$ to solve the above equation and obtain the interactive dynamic response $\mathcal{D}(t)$.
@@ 5 Dataset
To facilitate the learning of complex 3D tree dynamics, we construct 4DTree, a large-scale 4D tree dataset containing $8,786$ animated tree meshes. Each instance provides a $100$-frame physically plausible animation sequence that captures realistic and fine-grained wind-induced motion patterns across different tree shapes, sizes, and branching structures.
To create 4D tree data, a straightforward approach is to use commercial physics-based simulation software, but this is time-consuming and impractical for large-scale datasets. Instead, we adopt the method from [45], which models trees as hierarchical branching structures and simulates wind-induced motion by treating stems as elastic rods coupled through oscillators, implemented via the Sapling Tree Gen add-on in Blender. Still, this approach involves many sensitive parameters, which, if set improperly, can cause unstable oscillations or unrealistic shapes. To ensure quality and consistency of our dataset, we adopt a three-stage pipeline during production: parameter tuning, automatic validation with scripts, and final visual filtering by human reviewers. Through this process, we construct a clean and diverse 4D tree dataset with complex dynamics. For Detailed procedures and data samples, please refer to the supplementary material.
@@ 6 Experiments
Implementation. We train our model from scratch without pre-trained models, taking 3.5 days on 8 RTX 4090 GPUs with a batch size of 48. During the first 40,000 iterations, we train the model using only the $\mathcal{L}_{DM}$ loss. Then, we introduce the $\mathcal{L}_{LSS}$ loss and continue training for an additional 30,000 iterations. For the $\mathcal{L}_{LSS}$ loss, we use the 5 nearest neighbors of each point, and both $\alpha$ and $\lambda$ are set to 0.5. We set the resolution of the sparse voxel spectrum to $128^{3}$, with an input resolution of $512^{3}$ for the sparse voxel encoder. This results in voxel latent conditions of dimensionality $d=128$ at the $128^{3}$ resolution. When binding 3DGS primitives, we assign five Gaussians per face. We use the AdamW optimizer with an initial learning rate of $1\times 10^{-4}$, which is halved every 20,000 iterations. During inference, we employ DDIM [39] with 100 sampling steps.
Evaluation metrics. To evaluate our method in real-world scenarios, we collect a test set of 13 real-world trees. For evaluation metrics, we follow prior works [52, 6] and use CLIP ViT-B/32 [33] to measure both visual realism and temporal coherence. Specifically, we compute CLIP-I distance as the average CLIP distance between each frame and the input view, and CLIP-T distance as the average CLIP distance between consecutive frames. Furthermore, we conduct a user study on the rendered videos, focusing on four key aspects: motion authenticity (MA), motion complexity (MC), 3D structural consistency (3DSC), and visual quality (VQ). Below, we compare the results of motion generation and interaction simulation, respectively.
@@ 6.1 Comparison of 3D Animation
For 3D animation, we compare our method with the 4D generation approaches 4DGen [52] and SVD 2.0 [51]. As shown in Fig. 3, the results of 4DGen and SVD 2.0 often contain artifacts in fine structures. Moreover, when the tree structures become more complex, these baselines may fail to converge due to degraded motion generation in their underlying VDMs. For quantitative comparison, the results in Table 1 show that our method consistently outperforms these baselines across all metrics. Additional animation results are provided in the supplementary material, and we strongly recommend viewing the video results on our webpage for a clearer visual assessment.
@@ 6.2 Comparison of Interactive Simulation
For interactive dynamic simulation, we compare against state-of-the-art baselines PhysGaussian [48] and PhysFlow [26]. As shown in Fig. 4, PhysGaussian and PhysFlow often produce unrealistic plastic deformations. Specifically, PhysGaussian deforms slowly with little rebound, while PhysFlow exhibits partial recovery but still lacks fine-grained elasticity at the branch and leaf level, producing overly global responses. In contrast, our method produces natural and elastic motions, with branches and leaves exhibiting distinct behaviors. Quantitative results from four viewpoints and simulation time are reported in Table 1. Our method not only outperforms the baselines on most metrics but also significantly reduces the simulation time. For each frame, our method takes only about 18 ms for simulation, with 13 ms for mesh motion computation via modal analysis, 2.57 ms for Gaussian deformations calculation, and 2.65 ms for rendering, achieving a real-time interaction. Note that PhysFlow requires additional parameter optimization, which results in significantly longer runtime. More simulation results can be found in the supplementary material.
6.3 Comparison with prior physically-based tree animation work
We further conduct a comparative experiment against traditional tree animation methods. As the implementations of several existing methods are not publicly available, we choose [45] as a representative baseline. Note that a major difference from most prior work is that our method can be directly applied to real scanned 3D trees without heavy manual preprocessing, as our voxel-based representation is more robust to scanning errors. For fairness, however, we use the clean 3D tree models used in their algorithm rather than reconstructed trees with scanning artifacts. We perform a user study following the evaluation protocol in Sec. 6, and the results are shown in Table 2. The results show that even when using clean 3D inputs for comparison, our method consistently delivers superior visual quality across multiple evaluation metrics, demonstrating its effectiveness and robustness.
Table 2: Quantitative comparison with traditional tree animation methods.
[…]
Weber [45] 48.57% 37.14% 45.71% 34.29% 42.14%
Ours 51.43% 62.86% 54.29% 65.71% 57.86%
@@ 7 Limitation and Conclusion
7.1 Limitations
@@
7.2 Conclusion
In this paper, we present DynamicTree, a novel framework for animating 3DGS trees. By introducing the sparse voxel spectrum representation, our method enables efficient long-term real-scanned tree motion generation and real-time dynamic response to external forces. Furthermore, we also introduce a large-scale synthetic 4D tree dataset to support learning-based tree motion generation. Experimental results demonstrate that our approach achieves high-quality tree motion with strong temporal coherence and physical plausibility.
Appendix A Dataset
As discussed in the paper, we implement the method of [45] as a plugin in Blender. However, generating a large-scale 4D tree dataset using this method remains challenging, as it involves numerous parameters and random sampling often produces invalid or unrealistic 4D trees. To ensure data quality and consistency, we adopt a three-stage pipeline to construct our dataset:
Figure 6: Examples from our 4DTree dataset. For clarity of visualization, the leaves and trunk are rendered using two simplified material configurations.
@@ Appendix A Dataset
Appendix B Network
The sparse encoder and sparse voxel diffusion U-Net are adapted from the basic modules proposed in XCube to better fit our conditioning input and spectral output. We report key parameter settings of these two components in Table 4.
Appendix C Visualization results
Further visualizations and analytical details of the real-world 3D tree animations and interactive dynamic simulations are presented in Fig. 8 and Fig. 7. Moreover, to facilitate a comprehensive perceptual and qualitative evaluation of our method, we strongly recommend reviewing the video results on our webpage.
Figure 8: More results of real-tree 3D animation. For each scene, we visualize the space-time slices of depth and RGB videos from two viewpoints. We also show the scene flow of three mesh point cloud frames $(t_{1}=30,t_{2}=50,t_{3}=80)$ in the generated sequence, with color coding following the strategy used in [17], where the movements $(u,v)$ along the $x$ and $z$ directions are encoded using standard optical flow coloring.